OpenStudy (anonymous):

which equation represents y=-x^2+4x-1 in vertex form

OpenStudy (mathmale):

Welcome to OpenStudy! Glad you're here! You have a quadratic equation here. Your task is to re-write it so that the coordinates of the vertex of this parabola are obvious.

OpenStudy (mathmale):

The correct form of the equation of a quadratic for that purpose is\[y=a(x-h)^2+k\]

OpenStudy (mathmale):

So, basically, you need to re-write the given equation in that form. How would you do that?

OpenStudy (mathmale):

\[y=-x^2+4x-1 \rightarrow y=a(x-h)^2+k\]

OpenStudy (mathmale):

Hint: your " a " is going to be negative.

OpenStudy (mathmale):

@johnweldon1993: Thanks for your input, John. But wouldn't there be a negative coefficient, " a ", in this picture? Please, do your best to guide maddierawr through the solution of this problem without doing the actual work for her, OK? @maddierawr? Where are we? What's our goal? What's the next step towards identifying the coordinates of the vertex of this parabola?

OpenStudy (anonymous):

Just complete the square and you are done.

OpenStudy (anonymous):

\[ y=-x^2+4 x-1=-\left(x^2-4 x+4-4\right)-1=\\-\left(x^2-4 x+4\right)+4-1=-(x-2)^2 +3 \]

OpenStudy (mathmale):

Note how eliassaab has taken that negative sign out of the original expression, which is absolutely the correct thing to do. @maddierawr : You've gotten a lot of info here already; hope it's not overwhelming. Please let us know where you are at this point: what do you understand, and what do you need to know to finish solving this problem?